Chess Top uses the periodic inventory system. For the current month, the beginning inventory

consisted of 480 units that cost $65 each. During the month, the company made two purchases:

720 un

its at $68 each and 360 units at $70 each. Chess Top also sold 1,200 units during the

month. Using the average cost method, what is the amount of ending inventory

(round to the

nearest dollar)
Answer #1

Ending Inventory = 480+720+360 - 1200 = 360 units

Weighted Average Cost per unit = (480*65+720*68+360*70)/(480+720+360) = $67.54 per unit

Ending inventory = 360*67.54 = $24,314